3. A 20 m deep well with diameter 7 m is dug and the earth from digging is evenly spread out to form a platform 22 m by 14 m. Find the height of the platform.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
It is given that the shape of the well is in the shape of a cylinder with a diameter of 7 m So, radius = 7/2 m Also, Depth (h) = 20 m Volume of the earth dug out will be equal to the volume of the cylinder ∴ Volume of Cylinder = π×r2×h = 22×7×5 m3 Let the height of the platform = H Volume of soil from well (cylinder) = Volume of soil used to make such platform π×r2×h = Area of platform × Height of the platform We know that the dimension of the platform is = 22×14 So, Area of platform = 22×14 m2 ∴ π×r2×h = 22×14×H ⇒ H = 2.5 m